OpenStudy (anonymous):

Simplify. Express using positive exponents. *****PROBLEM REQUIRES EXPONENTS. READ COMMENTS FOR THE PROBLEM*****

OpenStudy (anonymous):

\[10^{15}\times 10^{4}\]

hero (hero):

In general, when multiplying exponents, the following applies: a^b x a^c = a^(b+c)

OpenStudy (anonymous):

I know. but would the answer be \[100^{19}\] or \[10^{19}\]

hero (hero):

The only thing you do is add the exponents. That's it.

OpenStudy (anonymous):

Ok! Thanks!!!